OpenStudy (anonymous):

how do you solve this Given that f(x) = x2 + 2x + 3 and g(x) = quantity of x plus four, over three, solve for f(g(x)) when x = 2. 2 5 11 33

OpenStudy (lynfran):

f(x)=x^2+2x+3 and g(x)=(x+4)/3 f(g(x)) = ((x+4)/3)^2+2((x+4)/3)+3 = (x+4)^2/9+((2x+8)/3) +3 =((x^2+8x+16)/9)+((2x+8)/3)+3 =(((2)^2+8(2)+16))/9)+((2(2)+8)/3)+3 =((4+16+16)/9))+((4+8)/3)+3 =??

OpenStudy (anonymous):

11?right

OpenStudy (lynfran):

correct

OpenStudy (anonymous):

alright thanks

OpenStudy (lynfran):

welcome